# Product-Stability: Provable Convergence for Gradient Descent on the Edge of Stability

### 03 Apr 2026 v1

@@ B.1 Phase I
$\gamma_{t+2}-\gamma_{t}$ $=O(\eta^{2}g_{t}^{2})$
@@
$\hat{Z}(z_{t+2})-\gamma_{t+2}=(1-\hat{Z}^{\prime}(\zeta_{t})\gamma_{t}l^{\prime\prime}(\theta_{t})g_{t})(\hat{Z}(z_{t})-\gamma_{t})+O(\eta^{2}g_{t}^{2})$
We split into two cases:
Case 1: $\hat{Z}(z_{t})+K\eta^{2}<\gamma_{t}$.
Indeed, looking at Equation 20, we know that $\gamma_{t}l^{\prime\prime}(\theta_{t})g_{t}(\hat{Z}(z_{t})-\gamma_{t})$ has the same sign as $z_{t}-z_{*}$. Moreover, using Lemma B.1, observe that $\hat{Z}^{\prime}(\zeta_{t})\gamma_{t}l^{\prime\prime}(\theta_{t})g_{t}=\Theta(g_{t}^{2})$, so we can choose $K$ sufficiently large such that there exists constant $C_{1}$ such that
@@
$(\hat{Z}(z_{0})-\gamma_{0})-(\hat{Z}(z_{T_{0}})-\gamma_{T_{0}})$ $=\sum(\hat{Z}(z_{t})-\gamma_{t})-(\hat{Z}(z_{t+2})-\gamma_{t+2})$
$\geq C_{1}\sum g_{t}^{2}(\hat{Z}(z_{t})-\gamma_{t})$
$\geq C_{1}K\eta^{2}\sum g_{t}^{2}$
On the other hand $\gamma_{t+2}-\gamma_{t}=O(\eta^{2}g_{t}^{2})$, so there exists $C_{2}$ such that
@@
$\leq\sum C_{2}\eta^{2}g_{t}^{2}$
$\leq\frac{C_{2}}{C_{1}K}[(\hat{Z}(z_{0})-\gamma_{0})-(\hat{Z}(z_{T_{0}})-\gamma_{T_{0}})]$
Choosing $K$ sufficiently large guarantees that this quantity is less than $\frac{\delta}{2}$, as desired.

### 13 Apr 2026 v2

@@ B.1 Phase I
$\gamma_{t+2}-\gamma_{t}$ $=-4\eta^{2}(g_{t+1}+g_{t})z_{t}+O(\eta^{2}g_{t}^{2})$ (21)
$=-4\eta^{2}l^{\prime\prime}(\theta_{t})g_{t}z_{t}(\hat{Z}(z_{t})-\gamma_{t})+O(\eta^{2}g_{t}^{2})$
@@
$\hat{Z}(z_{t+2})-\gamma_{t+2}=(1-\hat{Z}^{\prime}(\zeta_{t})\gamma_{t}l^{\prime\prime}(\theta_{t})g_{t}+4\eta^{2}l^{\prime\prime}(\theta_{t})g_{t}z_{t})(\hat{Z}(z_{t})-\gamma_{t})+O(\eta^{2}g_{t}^{2})$
We split into three cases:
Case 1: $\hat{Z}(z_{t})+K\eta^{2}<\gamma_{t},|g_{t}|\geq M\eta^{2}$.
Now, looking at Equation 20, we know that $\gamma_{t}l^{\prime\prime}(\theta_{t})g_{t}(\hat{Z}(z_{t})-\gamma_{t})$ has the same sign as $z_{t}-z_{*}$. Moreover, using Lemma B.1, observe that $\hat{Z}^{\prime}(\zeta_{t})\gamma_{t}l^{\prime\prime}(\theta_{t})g_{t}=\Theta(g_{t}^{2})$, so we can choose $K,M$ sufficiently large such that there exists constant $C_{1}$ such that
@@
$(\hat{Z}(z_{0})-\gamma_{0})-(\hat{Z}(z_{T_{0}})-\gamma_{T_{0}})$ $=\sum_{t}(\hat{Z}(z_{t})-\gamma_{t})-(\hat{Z}(z_{t+2})-\gamma_{t+2})$
$\geq C_{1}\sum_{t}g_{t}^{2}(\hat{Z}(z_{t})-\gamma_{t})$
$\geq C_{1}K\eta^{2}\sum_{t}g_{t}^{2}$
On the other hand, using Equation 21, there exists $C_{2}$ such that
@@
$\leq C_{2}\eta^{2}[(\hat{Z}(z_{0})-\gamma_{0})-(\hat{Z}(z_{T_{0}})-\gamma_{T_{0}})]+\sum C_{3}\eta^{2}g_{t}^{2}$
$\leq\left(C_{2}\eta^{2}+\frac{C_{3}}{C_{1}K}\right)[(\hat{Z}(z_{0})-\gamma_{0})-(\hat{Z}(z_{T_{0}})-\gamma_{T_{0}})]$
Choosing $K$ sufficiently large and $\eta$ sufficiently small guarantees that this quantity is less than $\frac{\delta}{2}$, as desired.
@@
Case 3: $|g_{t}|\leq M\eta^{2}$.
Since we are in the EoS regime, the update Equation 9 has a linearly unstable critical point at $z_{*}$. Hence the iterates diverge away from $z_{*}$ until we enter Case 1.